OpenStudy (anonymous):

Write a function for the graph described as a transformation of y=x^2. 12. y=x^2 experiences a vertical stretch of factor 2 and then a shift right of 3 units.

OpenStudy (anonymous):

\[y=a(x-h)^2+k\] a = vertical stretch factor h = horizontal shift k = vertical shift

OpenStudy (anonymous):

you don't have a vertical shift, so ignore the k = 0. Plug in the numbers for a and h.

OpenStudy (anonymous):

\[y=2(x-3)^{2}\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

can you check another question i have? @peachpi

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y=x^2 experience a shift left by 2 units, then a horizontal shrink of factor 1/2 then a shift down 5 units my answer : y=(x-1/2)^2

OpenStudy (anonymous):

horizontal shrinks/stretches are multiplied by the x from the inside. \[y=a(bx-h)^2+k\] Use that as your basic formula. a, h, and k are explained above. \(b\) is the horizontal shrink/stretch factor. The "default" numbers for a and b are 1, and for h and k it's 0. In your problem you have shift left, so h = -2 horizontal shrink, so b = 1/2 vertical shift down, so k = -5. They don't say anything about vertical stretch/shrink, so leave it as 1. \[y=\left( \frac{ 1 }{ 2 }x+2 \right)^2-5\]

OpenStudy (anonymous):

thanks for explaining

OpenStudy (anonymous):

you're welcome.

OpenStudy (anonymous):

fyi pictures are better than words :) http://www.regentsprep.org/regents/math/algtrig/atp9/funclesson1.htm